# Preserving Privacy in Large Language Models: A Survey on Current Threats and Solutions

### v1

@@ 2 Preliminaries
Large language models are deep learning systems trained on vast amounts of not-annotated text data to predict the likelihood of word sequences. Currently, most of the generative language models are trained in an autoregressive fashion, meaning they generate text by predicting each word sequentially based on the previous words in the sequence - usually from left to right. A model trained in this way is capable of generating and manipulating language effectively, which has been shown useful across a broad set of tasks and scenarios. However, this was not always the case. Before autoregressive models, there were masked language models (MLMs) like BERT (Devlin et al., 2019). MLMs are trained by masking certain words in a sentence and then predicting those masked words based on the surrounding context. They are much smaller than LLMs, with BERT large having 340 million parameters and they are also much less capable. Language modeling is not a new task: the groundwork for it was laid in 1948 by Claude Shannon (Shannon, 1948), and the task was extensively investigated for decades, but the recent advances with the introduction of the Transformer architecture (Vaswani et al., 2023) and the rollout of GPT-3 in 2022 (Brown et al., 2020b) shed new light on the task and this new class of very big language models, namely large language models. Training a Large Language Model from scratch, sometimes referred to as pre-training, is usually very computationally intensive given the scale of these models and requires several resources generally available only to a few organizations in the world (Sharir et al., 2020; Hadi et al., 2023). For example, to give a sense of this scale, GPT-4 is rumored to have around 1.8 trillion parameters with an estimated training cost of around \$63 million at the time of its release.
@@
With these models, a new training paradigm arose: the pre-training/fine-tuning paradigm. Pre-training is usually on the next token prediction task, while fine-tuning is on the downstream task (which can still be the next token prediction, but it can also be something else like a simple classification task). Organizations building these large language models typically pre-train them on extensive text corpora and then fine-tune them for specific tasks before making them accessible to the public. White-box access allows users to download the model’s weights and see or modify the model’s internal parameters and architecture, enabling detailed inspection and customization. In contrast, black-box access is usually offered through APIs or web interfaces, where users can input data and receive outputs without any visibility into the model’s internal workings or the ability to modify its parameters. Researchers and scientists can efficiently adapt the models to their specific tasks and needs through fine-tuning, which uses a relatively small fraction of the original pre-training data and iterations. Since these are language models, it is possible to interact with them in natural language, by feeding them sentences in input, which in this context are called prompts. This means that black box models can still be adapted by conditioning them with specific prompts, either in a zero-shot or few-shot fashion (Brown et al., 2020a).
@@ 2.1 Differential Privacy
$\delta$ is the parameter that allows for a small probability of failure of the privacy guarantee. Ideally, $\delta$ should be very small, close to 0, which indicates a lower chance of the algorithm failing to maintain privacy.
@@
Rényi Differential Privacy A randomized algorithm $\mathcal{M}:\mathcal{D}\mapsto\mathcal{R}$ is said to have $\epsilon$-Rényi differential privacy of order $\alpha$, or $(\alpha,\epsilon)$-RDP for short, if for any adjacent $D,D^{\prime}\in\mathcal{D}$ it holds that $D_{\alpha}\left(f(D)\|f\left(D^{\prime}\right)\right)\leq\epsilon$
@@
By providing a mathematically rigorous definition of privacy guarantees, Differential Privacy has become a key principle for designing privacy-preserving algorithms, especially in scenarios where personal data is involved. From this mathematical definition, some interesting properties arise. Among these, there are composition theorems that allow the computation of the exact privacy loss through many applications of the DP algorithm, which corresponds to the privacy budget. Sequential composition theorem (Kairouz et al., 2015) applies when multiple differentially private mechanisms are applied to the same dataset in sequence. The theorem quantifies how the privacy guarantees accumulate over the sequence of operations. Parallel composition theorem (McSherry, 2009) applies when multiple differentially private mechanisms are applied to disjoint subsets of the dataset. This theorem shows that the privacy guarantee of the overall process is governed by the worst-case privacy guarantee of any single mechanism. Differential Privacy also exhibits a Post-Processing property that states that once data has been processed through a DP mechanism, the output retains its privacy guarantees regardless of any further processing.
Differential privacy was born as a statistic analysis technique for databases. Initially, it was used to query databases without compromising the privacy of anybody enrolled, and that is why we discussed about queries and compounding privacy over them to compute the total privacy loss or privacy budget. Luckily, all this came in handy in recent years, with the worldwide diffusion of deep neural networks, when differential privacy was applied to them. That is when databases became datasets and queries became (mostly training) iterations. We will see more in detail what is the result of this transition in the next paragraph §2.1.
@@ 2.2 Deep Learning with Differential Privacy
Differential privacy was initially devised for database analysis, but it is possible to apply it to deep learning models in several ways. In practice, since we primarily work with deep learning models, we will often refer to datasets rather than databases. However, this distinction does not entail any practical difference in the context of our discussion. Figure 2 shows how Differential Privacy can be incorporated in 1) the input data, 2) the gradients, 3) the output data, 4) the labels or even in the loss function 5). The most used technique is by training using DP-SGD (Differentially Private Stochastic Gradient Descent), while DP-Adam is less used - we discuss the reasons for this in §5.1.2. Both of these techniques involve gradient manipulation, falling under the point 2) in Figure 2. Initially presented by Song et al. (2013), DP-SGD is a variation of standard Stochastic Gradient Descent (SGD) that enhances privacy by incorporating gradient clipping and adding Gaussian noise. In DP-SGD, gradients are first clipped to a maximum norm to limit the influence of any single data point. Then, Gaussian noise is added to the average of these clipped gradients, ensuring that the updates to the model parameters maintain differential privacy by obscuring the contribution of individual data points. It is possible to consider the several layers separately in order to have different clipping thresholds, and the amount of noise should be calibrated over the sensitivity of the function (in this case, the network), as explained by Dwork et al. (2006b). The sensitivity of a function f indicates how much the function’s output changes in response to changes in its input. In DP, there is also a focus on privacy costs, so DP-SGD needs a way to compute them. In Abadi et al. (2016), along with a refined version of DP-SGD, the authors present a novel technique for privacy accounting during training, which is called moments accountant. The composability property of DP allows to compute the privacy cost at every iteration and then sum them up. Privacy cost is computed in an asymptotic way, and while the mathematical details are beyond the scope of this survey, it is worth noting that moments accountant saves a non-negligible factor in the computation with respect to previous techniques, allowing to guarantee a lower privacy budget with the same magnitude of injected noise.
@@
Also gradient clipping is an important step in the DP learning process. Gradient clipping can be done in several ways, it is possible to use a per-sample clipping (Lee & Kifer, 2020) and this can even be done automatically (Bu et al., 2022a), or a flat clipping style (also called all-layer) (Abadi et al., 2016) or something in between, like group-wise clipping (McMahan et al., 2017). As stated by Bu et al. (2023), the first leans towards better memory accuracy while the second favours memory efficiency. Another option to enforce DP is PATE (Private Aggregation of Teacher Ensembles) (Papernot et al., 2016) and its variations (Papernot et al., 2018; Tian et al., 2022; Boenisch et al., 2022). In this method, several teacher models are trained on private data, and then a student model is trained on the predictions given by a majority vote with added noise made by the ensemble on public data. In this way, the actual model never sees the data, and the individual models (and, for extension, their data) cannot influence the outcome too much. A new, alternative way to enforce DP, as shown by Du et al. (2023b), is to perturb only forward pass embeddings, and according to the experimental evidence, it offers stronger performance with a lower vulnerability to attacks and more efficient memory usage. DP has been shown to work in several settings like NLU (Qu et al., 2021; Dupuy et al., 2021), image classification (De et al., 2022), diffusion models for image generation (Dockhorn et al., 2022; Ghalebikesabi et al., 2023), neural machine translation (Igamberdiev et al., 2023) and, of course, language modeling (Dinh & Fioretto, 2023) and LLMs in particular (Carranza et al., 2024). All these related works prove the interest of the scientific community in the problem and the commitment to finding an effective solution. Differentially private models are already in place with several solutions, and they work to different extents based on models, data and privacy budgets.
@@
Federated learning (McMahan et al., 2023) is a machine learning approach that enables model training across decentralized devices or servers holding local data samples without exchanging them. Instead of centralizing all data in one location, the model is trained collaboratively across multiple devices or servers. Each local device computes updates to the model based on its local data, and only these updates are shared and aggregated to improve the global model. Federated learning aims to preserve privacy and security by keeping sensitive data localized while still benefiting from the collective knowledge gained during the collaborative training process. It allows to be involved in the training of a model without fully trusting the external entity training the model (which can be a government or a company), because you are not giving up your actual data. As we will see more in detail later, this technique does not really manage to protect privacy by itself because gradient updates, even after aggregation, can be reversed to recover training data §3.3.2. A big advantage of federated learning is allowing to train better models thanks to the higher availability of high-quality local data, like the case of next word prediction for smartphone keyboards (Hard et al., 2018).
@@
Secure Multi-Party Computation (SMPC) (Yao, 1982; Zhao et al., 2019a) is a communication protocol that enables several parties to jointly compute a function over their private inputs, without revealing those inputs to each other. This way, each participant’s data remains confidential while still allowing for the collective computation of the function. It is particularly useful in scenarios where data privacy is crucial, yet there is a need for collaborative computation, such as in privacy-sensitive data analysis and secure voting systems. There are several ways to enforce SMPC, like secret sharing (Shamir, 1979; Blakley, 1899), homomorphic encryption (Kim & Yun, 2021; Joo & Yun, 2014), zero-knowledge proofs (Goldreich & Oren, 1994; Feige et al., 1988) and oblivious transfer (Chailloux et al., 2013; Cheong et al., 2009). While these techniques can enforce strong data security, they also add a considerable computational overhead that makes them difficult to apply in situations where the learning process is already heavy.
@@ 3 Privacy Attacks
As the training data becomes huge, it is increasingly urgent to understand whether models tend to expose the data on which they have been trained since it is more likely for them to capture sensitive information. A broad research area is devoted to understanding whether deep learning algorithms memorize training examples and, hence, can be used to reconstruct (potentially sensitive) training information. While some works question the effect of memorization during generation – meaning that texts generated starting from prompts extracted from training are different from the original ones (McCoy et al., 2023) – the discussion around memorization and its consequent risks is growing as larger models become more popular. In some cases, memorization is discussed as the key to good performance since data are often characterized by long-tailed distributions, hence, for a model, it is hard to distinguish between outliers and actually informative data points (Feldman & Zhang, 2020). However, memorization of training data by a model allows an attacker to reconstruct sensitive information or to understand whether a certain example was used during training. It has been shown by Carlini et al. (2023a) that it is possible to reconstruct training data with only black-box access.
For LLMs, it might be sufficient to input the right prompt, which is not necessarily an adversarial one (Carlini et al., 2021), in order to condition a model to generate sequences that are exactly memorized from the training data. Attacks that allow verbatim extraction of training data are generally referred to as Training Data Extraction attacks. When LLMs are fine-tuned with Reinforcement Learning with Human Feedback (RLHF), they tend to be more robust and avoid revealing the training set examples and, in particular, revealing sensitive information. However, some adversarial prompts have emerged to condition a model to show some unintended behaviour: since sensitive information is memorized by a model, RLHF is used to prevent the model from disclosing such information, but it does not ensure that new attacks cannot retrieve it. We discuss both non-adversarial and adversarial prompting in Section 3.1.
In some cases, an attacker may not only have access to a model but also to some data possibly included in the training phase of the target model. In this scenario, a Membership Inference Attack (MIA) can be performed: the attacker aims to claim that the information is included in the training set of a threatened model. The initial formulation of those attacks (Shokri et al., 2016) has also been extended against larger models and Transformer-based Language Models in particular. Since some sequences are memorized by a model, overfitting over them can be detected and used to attest to the presence of certain data in the original training data. In Section 3.2 we discuss those kind of attacks.
When private data owned by different holders need to be used to train a joint model on them, protocols like federated learning (Section 2.2) seem to be the appropriate solution: in this scenario, data are not distributed across multiple servers, only model snapshots and updates are shared. However, also model outputs and gradients can be used to infer the input that justifies the observed model behavior. The model outputs or gradients can be inverted to obtain the sequence that justifies them, allowing an attacker to reconstruct private data without them being explicitly shared. In Section 3.3) we discuss Model Inversion Attacks as a method to track back model input from the model’s output or gradients.

### v2

@@ 2.1 Large Language Models
Large language models are deep learning systems trained on vast amounts of not-annotated text data to predict the likelihood of word sequences. Currently, most of the generative language models are trained in an autoregressive fashion, meaning they generate text by predicting each word sequentially based on the previous words in the sequence - usually from left to right. A model trained in this way is capable of generating and manipulating language effectively, which has been shown useful across a broad set of tasks and scenarios (Zhao et al., 2023). However, this was not always the case. Other popular techniques train LMs by masking certain words in a sentence and then predicting those masked words based on the surrounding context, like in the case of BERT (Devlin et al., 2019). These models are typically much smaller and less capable than modern LLMs (for instance, BERT-large has 340 million parameters). Language modelling is not a new task: the groundwork for it was laid in 1948 by Claude Shannon (Shannon, 1948), and the task was extensively investigated for decades, but the recent advances with the introduction of the Transformer architecture (Vaswani et al., 2017) and the rollout of GPT-3 in 2022 (Brown et al., 2020) shed new light on the task and this new class of very big language models, namely large language models. Training a Large Language Model from scratch, sometimes referred to as pre-training, is usually very computationally intensive given the scale of these models and requires several resources generally available only to a few organizations in the world (Sharir et al., 2020; Hadi et al., 2023). For example, to give a sense of this scale, GPT-4 is rumoured to have around 1.8 trillion parameters with an estimated training cost of around \$63 million at the time of its release.
@@
With these models, a new training paradigm arose: the pre-training/fine-tuning paradigm. Pre-training is usually on the next token prediction task, while fine-tuning is on the downstream task (which can still be the next token prediction, but it can also be something else like a simple classification task). Organizations building these large language models typically pre-train them on extensive text corpora and then fine-tune them for specific tasks before making them accessible to the public. White-box access allows users to download the model’s weights and see or modify the model’s internal parameters and architecture, enabling detailed inspection and customization. In contrast, black-box access is usually offered through APIs or web interfaces, where users can input data and receive outputs without any visibility into the model’s internal workings or the ability to modify its parameters. There’s also a middle ground called grey-box access, which allows users a partial view and limited interaction with the model’s internal workings. Users can gain some insights into specific aspects of the model, such as the architecture or certain parameters, but do not have full access to download or modify the model’s weights and internal components. This type of access might enable customization and tuning through pre-defined interfaces or tools, offering more flexibility and control than black-box access but not as much transparency or openness as white-box access. Researchers and scientists can efficiently adapt the models to their specific tasks and needs through fine-tuning, which uses a relatively small fraction of the original pre-training data and iterations. Since these are language models, it is possible to interact with them in natural language by feeding them sentences in input, which in this context are called prompts. This means that black box models can still be adapted by conditioning them with specific prompts, either in a zero-shot or few-shot fashion (Brown et al., 2020).
@@ 2.2 Differential Privacy
$\delta$ is the parameter that allows for a small probability of failure of the privacy guarantee. Ideally, $\delta$ should be very small, close to 0, which indicates a lower chance of the algorithm failing to maintain privacy. When $\delta$ is 0, it is defined $\varepsilon$-differential privacy, which is the strongest kind of differential privacy.
@@
The simplest method to achieve differential privacy in a query is by adding random noise to its answer. The challenge lies in balancing the noise added to meet differential privacy standards without rendering the answer too noisy to be useful. To assist with this, foundational mechanisms have been developed in the field, detailing the type and amount of noise required. One such mechanism is the Laplace mechanism.
Laplace Mechanism According to the Laplace mechanism, for a function $f(x)$ which returns a real number, the following definition of $F(x)$ satisfies $\epsilon$-differential privacy: $F(x)=f(x)+\operatorname{Lap}\left(\frac{\Delta f}{\epsilon}\right)$ where $\Delta f$ is the sensitivity of $f$, and $\operatorname{Lap}(b)$ denotes sampling from the Laplace distribution with centre 0 and scale $b$.
Sensitivity $\Delta f$ refers to the maximum change in the output of a function or query when a single individual’s data in the input dataset is added or removed. It quantifies how much a query’s result can vary with the smallest possible change in the dataset, which is crucial in determining the amount of noise that needs to be added to ensure differential privacy. A function with high sensitivity requires more noise to protect individual privacy, while a function with low sensitivity requires less. Sensitivity is a key factor in calibrating the noise for mechanisms like the Laplace mechanism to ensure that individual data points do not significantly affect the output, thereby maintaining privacy.
Global Sensitivity For a function $f:\mathcal{D}\rightarrow\mathbb{R}$ mapping datasets $(\mathcal{D})$ to real numbers, the global sensitivity of $f$ is defined as follows: $GS(f)=\max_{x,x^{\prime}:d\left(x,x^{\prime}\right)<=1}\left|f(x)-f\left(x^{\prime}\right)\right|$
Here, $d(x,x^{\prime})$ denotes the distance between two datasets $x$ and $x^{\prime}$, with the notion that two datasets are neighbours if their distance is 1 or less. The way this distance is defined significantly impacts the privacy definition we achieve, and we’ll explore the dataset distance metric in more detail later.
Global sensitivity is defined such that for any two neighbouring datasets $x$ and $x^{\prime}$, the difference between $f(x)$ and $f(x^{\prime})$ does not exceed $GS(f)$. This is referred to as "global" sensitivity because it is independent of the specific dataset being queried, applying to any pair of neighbouring $x$ and $x^{\prime}$.
Based on the noise added, there are several mechanisms in the literature. The Gaussian mechanism is an alternative to the Laplace mechanism, which adds Gaussian noise instead of Laplacian noise.
Gaussian Mechanism The Gaussian mechanism does not satisfy pure $\epsilon$-differential privacy but does satisfy $(\epsilon,\delta)$-differential privacy. According to the Gaussian mechanism, for a function $f(x)$ which returns a number, the following definition of $F(x)$ satisfies $(\epsilon,\delta)$-differential privacy: $\begin{array}[]{r}F(x)=f(x)+\mathcal{N}(0,\sigma^{2})\\ \text{where }\sigma=\frac{\Delta f\cdot\sqrt{2\log(1.25/\delta)}}{\epsilon}\end{array}$ where $\Delta f$ is the sensitivity of $f$, and $\mathcal{N}(0,\sigma^{2})$ denotes sampling from the Gaussian (normal) distribution with mean 0 and variance $\sigma^{2}$. Note that here (and elsewhere in these notes), $\log$ denotes the natural logarithm.
The primary mechanisms we’ve discussed (Laplace and Gaussian) are designed for numerical outputs, adding noise directly to the results. But what if we need to provide an exact answer without injecting noise yet still maintain differential privacy? The exponential mechanism addresses this by enabling the selection of the "best" element from a set while ensuring differential privacy. The "best" element is identified by using a scoring function that assigns scores to each item within a given set of options. The exponential mechanism ensures differential privacy by approximately maximizing the score of the chosen element. To preserve privacy, it might sometimes select an element that doesn’t have the highest score.
Exponential Mechanism The exponential mechanism satisfies $\epsilon$-differential privacy: 1. A set $\mathcal{R}$ of possible outputs is selected 2. A scoring function is specified $u:\mathcal{D}\times\mathcal{R}\rightarrow\mathbb{R}$ with global sensitivity $\Delta u$ 3. The exponential mechanism outputs $r\in\mathcal{R}$ with probability proportional to: $\exp\left(\frac{\epsilon u(x,r)}{2\Delta u}\right)$
The main practical distinction between the exponential mechanism and previous mechanisms like the Laplace mechanism is that the exponential mechanism’s output is always an element from the specified set. This is particularly advantageous when choosing an item from a finite set, where a noisy result would be inappropriate.
@@
Rényi Differential Privacy A randomized algorithm $\mathcal{M}:\mathcal{D}\rightarrow\mathcal{R}$ is said to have $\epsilon$-Rényi differential privacy of order $\alpha$, or $(\alpha,\epsilon)$-RDP for short, if for any adjacent $D,D^{\prime}\in\mathcal{D}$ it holds that: $D_{\alpha}(\mathcal{M}(D)\|\mathcal{M}(D^{\prime}))\leq\epsilon$
@@
By providing a mathematically rigorous definition of privacy guarantees, Differential Privacy has become a key principle for designing privacy-preserving algorithms, especially in scenarios where personal data is involved. From this mathematical definition, some interesting properties arise. Among these, there are composition theorems that allow the computation of the exact privacy loss through many applications of the DP algorithm, which corresponds to the privacy budget. Sequential composition theorem (Kairouz et al., 2015) applies when multiple differentially private mechanisms are applied to the same dataset in sequence. The theorem quantifies how the privacy guarantees accumulate over the sequence of operations. Parallel composition theorem (McSherry, 2009) applies when multiple differentially private mechanisms are applied to disjoint subsets of the dataset. This theorem shows that the worst-case privacy guarantee of any single mechanism governs the privacy guarantee of the overall process. Differential Privacy also exhibits a Post-Processing property that states that once data has been processed through a DP mechanism, the output retains its privacy guarantees regardless of any further processing.
Differential privacy was born as a statistic analysis technique for databases. Initially, it was used to query databases without compromising the privacy of anybody enrolled, and that is why we discussed queries and compounding privacy over them to compute the total privacy loss or privacy budget. Luckily, all this came in handy in recent years, with the worldwide diffusion of deep neural networks, when differential privacy was applied to them. That is when databases became datasets and queries became (mostly training) iterations. We will see more in detail what is the result of this transition in the next paragraph §2.2.
@@ 2.3 Deep Learning with Differential Privacy
Differential privacy was initially devised for database analysis, but it is possible to apply it to deep learning models in several ways. In practice, since we primarily work with deep learning models, we will often refer to datasets rather than databases. However, this distinction does not entail any practical difference in the context of our discussion. Figure 2 shows how Differential Privacy can be incorporated in 1) the input data, 2) the gradients, 3) the output data, 4) the labels or even in the loss function 5). The most used technique is by training using DP-SGD (Differentially Private Stochastic Gradient Descent), while DP-Adam is less used - we discuss the reasons for this in §5.1.2. Both of these techniques involve gradient manipulation, falling under point 2) in Figure 2. Initially presented by Song et al. (2013), DP-SGD is a variation of standard Stochastic Gradient Descent (SGD) that enhances privacy by incorporating gradient clipping and adding Gaussian noise. In DP-SGD, gradients are first clipped to a maximum norm to limit the influence of any single data point. Then, Gaussian noise is added to the average of these clipped gradients, ensuring that the updates to the model parameters maintain differential privacy by obscuring the contribution of individual data points. It is possible to consider the several layers separately in order to have different clipping thresholds, and the amount of noise should be calibrated over the sensitivity of the function (in this case, the network), as explained by Dwork et al. (2006b). The sensitivity of a function f indicates how much the function’s output changes in response to changes in its input. In DP, there is also a focus on privacy costs, so DP-SGD needs a way to compute them. In Abadi et al. (2016), along with a refined version of DP-SGD, the authors present a novel technique for privacy accounting during training, which is called moments accountant. The composability property of DP allows us to compute the privacy cost at every iteration and then sum them up. Privacy cost is computed in an asymptotic way, and while the mathematical details are beyond the scope of this survey, it is worth noting that moments accountant saves a non-negligible factor in the computation with respect to previous techniques, allowing to guarantee a lower privacy budget with the same magnitude of injected noise.
@@
Also gradient clipping is an important step in the DP learning process. Gradient clipping can be done in several ways. It is possible to use a per-sample clipping (Lee & Kifer, 2021) and this can even be done automatically (Bu et al., 2023a), or a flat clipping style (also called all-layer) (Abadi et al., 2016) or something in between, like group-wise clipping (McMahan et al., 2017b). As stated by Bu et al. (2024), the first leans towards better memory accuracy while the second favours memory efficiency. Another option to enforce DP is PATE (Private Aggregation of Teacher Ensembles) (Papernot et al., 2017) and its variations (Papernot et al., 2018; Tian et al., 2022). In this method, several teacher models are trained on private data, and then a student model is trained on the predictions given by a majority vote with added noise made by the ensemble on public data. In this way, the actual model never sees the data, and the individual models (and, for extension, their data) cannot influence the outcome too much. A new, alternative way to enforce DP, as shown by Du et al. (2023b), is to perturb only forward pass embeddings, and according to the experimental evidence, it offers stronger performance with a lower vulnerability to attacks and more efficient memory usage. DP has been shown to work in several settings like NLU (Qu et al., 2021; Dupuy et al., 2022), image classification (De et al., 2022), diffusion models for image generation (Dockhorn et al., 2023; Ghalebikesabi et al., 2023), neural machine translation (Igamberdiev et al., 2024) and, of course, language modelling (Dinh & Fioretto, 2023) and LLMs in particular (Carranza et al., 2024). All these related works prove the interest of the scientific community in the problem and the commitment to finding an effective solution. Differentially private models are already in place with several solutions, and they work to different extents based on models, data and privacy budgets.
@@
Federated learning (McMahan et al., 2017a) is a machine learning approach that enables model training across decentralized devices or servers holding local data samples without exchanging them. Instead of centralizing all data in one location, the model is trained collaboratively across multiple devices or servers. Each local device computes updates to the model based on its local data, and only these updates are shared and aggregated to improve the global model. Federated learning aims to preserve privacy and security by keeping sensitive data localized while still benefiting from the collective knowledge gained during the collaborative training process. It allows to be involved in the training of a model without fully trusting the external entity training the model (which can be a government or a company), because you are not giving up your actual data. As we will see more in detail later, this technique does not really manage to protect privacy by itself because gradient updates, even after aggregation, can be reversed to recover training data §3.3.2. A big advantage of federated learning is allowing to train better models thanks to the higher availability of high-quality local data, like the case of next word prediction for smartphone keyboards (Hard et al., 2018).
@@
Secure Multi-Party Computation (SMPC) (Yao, 1982; Zhao et al., 2019a) is a communication protocol that enables several parties to jointly compute a function over their private inputs, without revealing those inputs to each other. This way, each participant’s data remains confidential while still allowing for the collective computation of the function. It is particularly useful in scenarios where data privacy is crucial, yet there is a need for collaborative computation, such as in privacy-sensitive data analysis and secure voting systems. There are several ways to enforce SMPC, like secret sharing (Shamir, 1979; Blakley, 1899), homomorphic encryption (Gentry, 2009; Kim & Yun, 2021; Joo & Yun, 2014), zero-knowledge proofs (Goldreich & Oren, 1994; Feige et al., 1988) and oblivious transfer (Rabin, 2005; Even et al., 1985; Brassard et al., 1986; Chailloux et al., 2013; Cheong et al., 2009). Homomorphic encryption is particularly interesting in this context as it allows the computation of a function on encrypted data. This makes it possible to perform training and inference of machine learning models on encrypted data. While these techniques can enforce strong data security, they also add a considerable computational overhead that makes them difficult to apply in situations where the learning process is already heavy.
@@ 3 Privacy Attacks
As the training data becomes huge, it is increasingly urgent to understand whether models tend to expose the data on which they have been trained since they are more likely to capture sensitive information. Moreover, the leakage of the original training data is not the only possible threat. In fact, the versatility of LLM is enhanced by the use of in-context demonstrations of the task that a user wants to perform (Brown et al., 2020). While being a useful technique, unconscious use can cause sensitive information – that should not be disclosed – to be included in these examples in context as well. For commercial systems, those data can also be leaked by exploiting vulnerabilities of the systems storing them (OpenAI, 2023). Even cyber security measures are insufficient since those data can be used in further training steps and subsequently leaked by a future version of the model. This aspect is also concerning for data with commercial value from a business perspective (Coles, 2023), and companies have decided, in some cases, to warn their employees about the responsible use of LLMs (Kim, 2023).
A broad research area is devoted to understanding whether deep learning algorithms memorize training examples and, hence, can be used to reconstruct (potentially sensitive) training information. While some works question the effect of memorization during generation – meaning that texts generated starting from prompts extracted from training are different from the original ones (McCoy et al., 2023) – the discussion around memorization and its consequent risks is growing as larger models become more popular. In some cases, memorization is discussed as the key to good performance since long-tailed distributions often characterize data. Hence, for a model, it is hard to distinguish between outliers and actually informative data points (Feldman & Zhang, 2020). However, memorization of training data by a model allows an attacker to reconstruct sensitive information or to understand whether a certain example was used during training. It has been shown by Carlini et al. (2023a) that it is possible to reconstruct training data with only black-box access.
For LLMs, it might be sufficient to input the right prompt, which is not necessarily an adversarial one (Carlini et al., 2021), in order to condition a model to generate sequences that are exactly memorized from the training data. Attacks that allow verbatim extraction of training data are generally referred to as Training Data Extraction attacks. When LLMs are fine-tuned with Reinforcement Learning with Human Feedback (RLHF), they tend to be more robust and avoid revealing the training set examples and, in particular, revealing sensitive information. However, some adversarial prompts have emerged to condition a model to show some unintended behaviour. Since sensitive information is memorized by a model, RLHF is used to prevent the model from disclosing such information, but it does not ensure that new attacks cannot retrieve it. We discuss both non-adversarial and adversarial prompting in §3.1.
In some cases, an attacker may not only have access to a model but also to some data possibly included in the training phase of the target model. In this scenario, a Membership Inference Attack (MIA) can be performed: the attacker aims to claim that the information is included in the training set of a threatened model. The initial formulation of those attacks (Shokri et al., 2017) has also been extended against larger models and Transformer-based Language Models in particular. Since some sequences are memorized by a model, overfitting over them can be detected and used to attest to the presence of certain data in the original training data. The effectiveness of MIA is also discussed in light of the training practices generally used to develop LLMs, such as training for a single epoch on a huge amount of data, which may influence the measure of overfitting used in these attacks. In §3.2, we discuss that kind of attack.
When private data owned by different holders need to be used to train a joint model on them, protocols like federated learning (§2.3) seem to be the appropriate solution: in this scenario, data are not distributed across multiple servers, only model snapshots and updates are shared. However, model outputs and gradients can also be used to infer the input that justifies the observed model behaviour. The model outputs or gradients can be inverted to obtain the sequence that justifies them, allowing an attacker to reconstruct private data without them being explicitly shared. We discuss Model Inversion Attacks as a method to track back model input from the model’s output or gradients in §3.3). Moreover, from those rather opaque sources, like model outputs, an attacker can gain information regarding a model’s parameters: those attacks, called Model Stealing attacks, while still not able to reconstruct entire LLMs, can potentially be directed against API-gated models, making them less opaque and, hence, more vulnerable against privacy attacks.
Finally, we notice that since fine-tuning is more expensive than in-context learning, many systems use in-context demonstration to specialize an LLM on specific tasks. Despite the flexibility of this method, some risks may emerge from using private information for in-context demonstrations. For example, a system that answers a user query by leveraging a prompted LLM could be vulnerable to privacy attacks performed by a malicious user (see §3.4 for further details.)